OpenStudy (anonymous):

Find the formula for the inverse function F(X)=1+Radicle(2+3x).

OpenStudy (anonymous):

I don't quite have an idea of how to solve this problem.Thanks for any suggestions.

OpenStudy (anonymous):

Start by writing the function like this : y = sqrt(2+3x)

OpenStudy (anonymous):

Then switch the variables, this way you'll have x= 1+ sqrt(2+3x)

OpenStudy (anonymous):

Oh, Okay I see what you did there! F(x) is just another way of saying Y equals. However, once the variable have been switched, what exactly is the next step? Would you now Subtract one from both sides to yield X-1=sqrt(2+ 3x)?

OpenStudy (anonymous):

Exactly! Sorry that I left you hanging. So your answer would be 1/3 (x^2-2x-1)=y

OpenStudy (anonymous):

Oh my, no worries! Thank you and so sorry I haven't replied I've just been getting all of my h.w. done! but again thank you so very much I really appreciate your help!! have a lovely night!